# Kinetic formulation of compartmental epidemic models

### v1

@@ 5 Examples
In the second example, in Subsec. 5.2, we consider a situation that unfortunately happens in several diseases when in certain localities, ranging from small villages to entire countries, large parts of the population behave as if the disease did not exist. Frequently, this is driven by misinformation campaigns. However, outside this region, susceptible individuals try to avoid infectious individuals, as is natural to expect.
Finally, Subsec. 5.3 considers that infectious individuals voluntarily confine themselves, represented by a decrease in the velocity of I individuals with respect to S and R individuals. In the previous discussion, the set of admissible velocities $V$ was the same for all classes of individuals, and considering a different set of admissible individuals introduces serious technical difficulties when dealing with the class-transition kernels $Q^{(T)}$. Therefore, we will consider this situation only in the one-dimensional case, in which formal calculations are significantly simpler.
@@ 5.1 The constant parameters case
Assume that $\alpha,\beta,\gamma,\eta,\xi$ are positive constants. In this case, $\kappa=vF(|v|)/\lambda$, and $\Theta=(\eta+\xi)vF(|v|)/\lambda$, and, therefore the coefficients of the mesoscopic equations are given by
@@
If we assume further that the space of admissible velocities is a circle or a sphere of radius $v_{0}$, for a movement in the two- or three-dimensional space, respectively, then $D=v_{0}^{2}/\lambda$ and $\Gamma=(\eta+\xi)v_{0}^{2}/\lambda$.
@@ 5.3 Confining infectious individuals
We will assume three set {±$\{\pm v_{S}\}$, {±$\{\pm v_{I}\}$, {±$\{\pm v_{R}\}$, $v_{j}>0$, representing possible velocities for S, I, and R individuals, respectively. Density of individuals will be given by $f_{j}^{\pm}=f_{j}(t,x,\pm v_{j})$, with the superscript sign indicating if the corresponding individuals move to the right or to the left. Mesoscopic densities are given by $\rho_{j}=f_{j}^{-}+f_{j}^{+}$.
@@
$0=Q^{(R)}_{j}[\mathbf{f}]=\lambda_{j}\left(\frac{f_{j,0}^{+}+f_{j,0}^{-}}{2}-f_{j,0}^{\pm}\right)\ ,$
@@
$\lambda_{j}\left(\frac{\kappa^{+}_{j}-\kappa^{-}_{j}}{2}\right)=Q^{(R)}_{j}[\kappa_{j}]=-\frac{v_{j}}{2}\ ,$
@@
$\Gamma_{j}=\frac{(\eta+\xi)v_{j}}{\lambda_{j}}\delta_{jS}\ .$
@@
$\partial_{t}\rho_{S}-\nabla_{x}\cdot\left(\frac{(\eta+\xi)v_{S}^{2}}{\lambda_{S}}(\nabla_{x}\rho_{I})\rho_{S}\right)-\nabla_{x}\cdot\left(\frac{v_{S}^{2}}{\lambda_{S}}\nabla_{x}\rho_{S}\right)=Q^{(T)}_{S}\left[\frac{\bm{\rho}}{2}\right]\ ,$
$\partial_{t}\rho_{j}-\nabla_{x}\left(\frac{v_{j}^{2}}{\lambda_{j}}\nabla_{x}\rho_{j}\right)=Q^{(T)}_{j}\left[\frac{\bm{\rho}}{2}\right]\ ,\qquad j=I,R\ ,$
@@ 6 Conclusions and future work
The idea of this work is to provide a model compatible with the well-known SIRS model in mathematical epidemiology, but with more flexibility to include several effects that cannot be easily implemented in spatially-homogeneous compartmental models.
We considered that individuals move freely in a bounded domain and contagion occurs when susceptible and infectious individuals are walking side by side, i.e., with the same velocity and at the same point in space. We used recent ideas both in kinetic models and in game theory to provide a model that is compatible, at least at the formal level, with the SIRS system of ordinary differential equations.
An important point missing from the present work is to show the links between the several different descriptive scales. More precisely, one important result to be discussed in a future work is the convergence of the solutions of the kinetic model given by Eq. (3) to the solution of the mesoscopic drift-diffusion-advection model given by Eqs. (9), (10), and (11). Mathematically, this means to study in which space, for which parameters, and for which initial conditions can we guarantee $\lim_{\varepsilon\to 0}\int_{V}f_{j,\varepsilon}\mathrm{d}v=\rho_{j}$. In the same vein, we should study how small perturbations of homogeneous initial conditions behave in time, i.e., to discuss if the solution of the SIRS system of ordinary differential equations is a good approximation of the more realistic non-homogenous problem, both in the kinetic and in the mesoscopic (drift-diffusion-advection) level, when the problem is close to the homogenous one. In particular, several important parameters, such as the basic reproductive number $\mathcal{R}_{0}$, are defined in the homogeneous case. As there is no universally accepted formulation of a non-homogeneous compartmental model, it is unclear how robust these concepts are in providing reliable information when dealing with real data, which is naturally non-homogeneous. In [WZ12], a generalization of $\mathcal{R}_{0}$ was provided for non-homogeneous epidemic models at the PDE level. Yet, we are not aware of the definition of any similar concept at the kinetic level.
Another missing point of the present work is to consider in more detail the global existence and uniqueness of solutions when the preferred direction kernel is non-zero. One possible approach, that goes in the direction of [CMPS04, Model 2] is to replace $v\cdot\nabla_{x}f_{I}$ by $\frac{1}{\varepsilon}\left(f_{I}(x+\varepsilon v)-f_{I}(x)\right)$, with an appropriate high-order correction. In principle, this should not change the macroscopic limit, but it may be more manageable for rigorous proofs.
Assume that the preferred direction kernel is of one of the two forms discussed in the previous paragraph. From the modeling point of view, we expect $\eta,\xi>0$ to have a regularizing effect on the solution of the kinetic model. More precisely, assume that solution $\mathbf{f}$ exists in a given space for $\eta=\xi=0$. Considering that I individuals are generated by S individuals, and for $\eta,\xi>0$, S individuals tend to avoid high concentrations of I individuals, and considering that the absence of S individuals close to the high concentration of I will provide a limitation in the emergence of new I individuals, we expect that a solution would also exist in this new situation.
The situation would be completely different if we allowed $\eta$ and/or $\xi$ to be negative. For the problem we are interested in, this would not be realistic; however, in the more general game-theoretical situation discussed in Subsec. 3.3, that could be of interest, modeling, for example, the formation of aggregates starting from a small perturbation of a spatially-homogeneous solution. This idea is close to the original chemotaxis problem [KS70], which inspired the kinetic model discussed in [CMPS04], one of the main sources of the present work. The general kinetic model with game-theoretical class transition kernels approach could be of interest in the study of pattern formation mediated by games played between cells, cf. [Tom97, DCS+09] for applications of evolutionary game theory in cancer modeling, or in the evolution of cooperation between rational agents, cf. [AH81] for the initial studies in the evolution of cooperation and [NBM94] for its extension to spacial games.
Another limitation of the model is that individual behavior is purely mechanistic. However, typical human behavior considers at the same time local and global information, a level of detail that cannot be explored in ODE models. Furthermore, this information can be contradictory. An explicit example will help to understand that point. Consider a new class of individuals, the Vaccinated one, and assume that vaccination is voluntary. Human behavior depends on global information $\int_{\Omega\times V}f_{I}\mathrm{d}v\mathrm{d}x$ (i.e., the endemic situation in the entire population) and in the behavior of close friends and relatives (local information, i.e., $f_{V}$, the new class of type V individuals), and could also depend on the immediate perception, which is also a local information. These values can be contradictory, i.e., even if a large outbreak is happening, a cluster of unvaccinated individuals can exist, especially in isolated communities that were not yet reached by the outbreak. It is not clear that simple modifications of the kinetic or the PDE model discussed is the right model to capture that feature, but it is clear that an ODE model will not, as there is no meaningful difference between local and global in the last case. Human behavior and vaccination is studied from the mathematical point of view since the seminal work by [BE04]; see also [DRSC16, CDPdCS24].
In the near future, we hope to provide some answers for at least some of the questions above.

### v2

@@ 5 Examples and applications
The third example, in Subsec. 5.3, considers a situation that unfortunately occurs in several diseases, whereby large parts of the population in certain localities, ranging from small villages to entire countries, behave as if the disease did not exist. This is frequently driven by misinformation campaigns. However, outside this region, susceptible individuals naturally try to avoid infectious individuals.
Finally, Subsec. 5.4 considers infectious individuals voluntarily confining themselves, which is represented by a decrease in the velocity of I individuals with respect to S and R individuals, as discussed at the beginning of this section. We demonstrate that confining I individuals reduces the intensity of disease spread.
5.1 A numerical exploration
Consider a large number of particles moving within a finite-size square lattice with periodic boundary conditions. Individuals are of type S, I, and R. $S(x,v)$ is the number of individuals of type S at $x$ moving at velocity $v$, with similar notation for $I(x,v)$ and $R(x,v)$.
Initially, all individuals are uniformly distributed with random initial velocities in a central square whose side is one-third the size of the full lattice. The majority are of type S, with a small number of type I individuals. No R individuals are initially present.
The update consists of two sequential steps: first, the moving direction is updated and the particle moves accordingly; then, class transitions are considered.
We start by moving each particle. For each individual, we consider two components of movement: the first in the horizontal direction and the second in the vertical direction. These are updated independently, resulting in eight admissible velocities corresponding to the set V: north, northwest, west, southwest, south, southeast, east and northeast. Consider an individual moving in one of the two main directions: horizontal or vertical. If the number of individuals moving in the same direction exceeds the average, there is a probability of changing their moving direction, denoted by a value of $\lambda\in(0,1]$. Otherwise, the direction remains unchanged. After performing the calculations for the horizontal and vertical directions independently, we update that individual’s moving direction. For example, if an individual at position $x$ with velocity $v$ moves northeast ($v^{\prime}=(1,1)$), then we decrease $I(x,v)$ by one and increase $I(x+(1,1),(1,1))$ by one. All updates are performed only after running the update rule for all particles (synchronous update).
We did not consider preferred direction kernels, i.e. we set all of the $\mathcal{Q}^{(P)}_{j}$ kernels to zero.
Once all the movements have been performed, the class transitions are considered. The parameters $\alpha$, $\beta$, and $\gamma$ are fixed. For each individual, a random number $r$ is drawn uniformly between 0 and 1. If the individual is of class S and $r_{1}<\beta I(x,v)/(S(x,v)+I(x,v)+R(x,v))$, then $S(x,v)$ decreases by one and $I(x,v)$ increases by one, indicating the S to I transition. If the individual is of class I and $r<\gamma$, then $I(x,v)$ decreases by one and $R(x,v)$ increases by one. Finally, if the individual is of type R and $r<\alpha$, then $R(x,v)$ decreases by one and $S(x,v)$ increases by one. Again, the update is synchronous.
Then, we repeat the movement and transition steps until the end of the simulation. We finish the simulation before the boundary conditions become relevant. See Fig. 3 for the evolution of the mesoscopic quantities, $\rho_{j}$.
Figure 3: Representation of the mesoscopic density $\rho_{S}$ (blue, left), $\rho_{I}$ (red, middle), and $\rho_{R}$ (right, green). More intense colors indicate higher concentrations. We consider the initial condition given by S and I individuals in the center of a square lattice of size $50\times 50$, with random initial velocities. Initial conditions are far from the boundary which are not reached for times considered. Initially, there are 10000 S individuals and 50 I individuals. We plot the configuration after 3, 6, 9, 12, and 15 steps, each step including the update of all particles. Class-transition parameters were given by $\alpha=0$ (i.e., immunity is permanent), $\beta=0.75$ (each encounter between S and I individuals has a probability of $\nicefrac{{3}}{{4}}$ to generate a new I individual), and $\gamma=0.5$ (i.e., in each time step, each I individual has a probability $\nicefrac{{1}}{{2}}$ to became R. The regression toward the mean parameter is $\lambda=0.5$, and we did not consider preferred directions, i.e., $\mathcal{Q}^{(P)}_{j}\equiv 0$.
@@ 5.2 The constant parameters case
Assume that $\alpha,\beta,\gamma,\eta,\xi$ are positive constants. In this case, $\kappa_{j}=vF(|v|)/\lambda$, and $\Theta=(\eta+\xi)vF(|v|)/\lambda$, see Eqs. (8) and (9). The coefficients of the mesoscopic equations are given by
@@
If we assume further that the space of admissible velocities is a hypersphere of radius $v_{0}$, for a movement in the two- or three-dimensional space, respectively, then $D=v_{0}^{2}/\lambda$ and $\Gamma=(\eta+\xi)v_{0}^{2}/\lambda$.
@@ 5.4 Confining infectious individuals
Given positive values $v_{S},v_{I},v_{R}>0$, we define three sets $V_{S}:=\{\pm v_{S}\}$, $V_{I}:=\{\pm v_{I}\}$, $V_{R}:=\{\pm v_{R}\}$, representing the set of possible velocities for S, I, and R individuals, respectively. Density of individuals will be given by $f_{j}^{\pm}=f_{j}(t,x,\pm v_{j})$, with the superscript sign indicating if the corresponding individuals move to the right or to the left. Mesoscopic densities are given by $\rho_{j}=f_{j}^{-}+f_{j}^{+}$.
@@
$0=\mathcal{Q}^{(R)}_{j}[\mathbf{f}]=\lambda_{j}\left(\frac{f_{j,0}^{+}+f_{j,0}^{-}}{2}-f_{j,0}^{\pm}\right)\ ,$
@@
$\lambda_{j}\left(\frac{\kappa^{+}_{j}-\kappa^{-}_{j}}{2}\right)=\mathcal{Q}^{(R)}_{j}[\kappa_{j}]=-\frac{v_{j}}{2}\ ,$
@@
$\Gamma_{j}=\frac{(\eta+\xi)v_{j}^{2}}{\lambda_{j}}\delta_{jS}\ .$
@@
$\partial_{t}\rho_{S}-\nabla_{x}\cdot\left(\frac{(\eta+\xi)v_{S}^{2}}{\lambda_{S}}(\nabla_{x}\rho_{I})\rho_{S}\right)-\nabla_{x}\cdot\left(\frac{v_{S}^{2}}{\lambda_{S}}\nabla_{x}\rho_{S}\right)=\mathcal{Q}^{(T)}_{S}\left[\frac{\bm{\rho}}{2}\right]\ ,$
$\partial_{t}\rho_{j}-\nabla_{x}\left(\frac{v_{j}^{2}}{\lambda_{j}}\nabla_{x}\rho_{j}\right)=\mathcal{Q}^{(T)}_{j}\left[\frac{\bm{\rho}}{2}\right]\ ,\qquad j=I,R\ ,$
@@ 6 Conclusions and future work
We begin this work by introducing a kinetic model to describe the pathogen dynamics within a population. We considered a population of individuals that move freely in a bounded domain. Contagion occurs when susceptible and infectious individuals are in close proximity for an extended period of time. In other words, pathogen transmission only occurs if an S and an I individual are at the same point of the domain and moving with the same velocity. We started by showing that the kinetic model is formally compatible SIRS system of ordinary differential equations, i.e., the former simplifies to the latter when dealing with solutions depending only on time.
The main advantage of using kinetic models is that they allow spatial phenomena, such as individual motility and spatial heterogeneity, to be included in the modelling. However, there are significantly more possibilities, and the present work has only begun to explore them.
One interesting application to be explored in future work is the introduction of vaccination, specifically allowing each individual in the population to freely choose whether to be vaccinated. The interaction between human behaviour with respect to vaccination and pathogen dynamics has been studied mathematically since the seminal work of [BE04]; see also [DRSC16, CDPdCS24]. Unlike models based on ordinary differential equations, spatially inhomogeneous models allow us to consider local and global information simultaneously, both of which are relevant to each individual’s decision on the central question: ”To be or not to be vaccinated?” Furthermore, local and global information can produce conflicting answers to this question. An explicit example will help to illustrate this point. Consider a new class of individuals: the vaccinated ones. Assume that vaccination is voluntary. In a minimalistic model, human behaviour depends on global information, i.e. the endemic situation in the entire population, and local information given by the densities $f_{j}$, where $j=S,I,R,V$. Clearly, at least transiently, an outbreak in one region can coexist with a disease-free region. Kinetic models in epidemiology may provide a richer description of reality by considering the attenuation of information due to distance and the use of non-physical variables, e.g. describing individual risk acceptance.
A second approach is to explore more general transition kernels. In the current work, the transition kernel was specifically designed to correspond to the SIRS model at a macroscopic level. However, any system of ordinary differential equations in the form of the replicator system can be obtained as the macroscopic limit of a suitably defined class transition kernel. The replicator system of differential equations is a key model in evolutionary game theory (see, for example, [HS98]). In this work, we built upon a recent reformulation of the SIRS model in replicator equation form (cf. [HC24]). To the best of our knowledge, the full relationship between kinetic models and game theory remains unexplored; however, see [AFMS21].
An important classical topic in the literature of kinetic models for chemotaxis is the existence of finite-time blow up versus global-in-time solution; in particular, a finite-time blow up may occur in the partial differential equation regime (mesoscopic level, in our parlance), i.e., in the Keller-Segel system of partial differential equations [KS70]. However, a formally compatible kinetic model may exhibit global existence [DP04, CMPS04]. This question, which is extremely interesting for certain kernels inspired by game theory, is a current topic of interest for future research. Let us be more specific.
In the present work, we considered that S individuals turn into I by proximity, and that S tend to escape from higher concentrations of I individuals, or are neutral to higher concentrations of I individuals. We proved existence and uniqueness of solutions of the kinetic model in the case $\mathcal{Q}^{(P)}_{j}\equiv 0$, i.e. $\eta,\xi=0$, however, we expect that $\eta,\xi>0$ will have a regularising effect on the solution of the kinetic model, and Thm. 2 will still hold in this case. The situation would be completely different if we allowed $\eta$ and/or $\xi$ to be negative, i.e., if individuals of class X change to Y and are attracted by high concentrations of Y individuals. Assume an evolutionary dynamics in which individuals with higher payoff reproduce at a higher rate and a rational behaviour in which they are attracted by environments in which they can expect an increase in payoff. These assumptions are natural in evolutionary game theory. In games of dominance or coordination games, one type will fixate while the other will become extinct [HS98]. In these cases, we expect a richer picture in which not only would finite-time blow up be possible, but also pattern formation, particularly in coordination games, as both types may fixate in different regions. The final outcome will possibly depend on the initial condition and deserves investigation. The general kinetic model with game-theoretical class transition kernels approach could be of interest in the study of pattern formation mediated by games played between cells, cf. [Tom97, DCS+09] for applications of evolutionary game theory in cancer modelling, or on the evolution of cooperation between rational agents, cf. [AH81] for the initial studies on that topic and [NBM94] for its extension to spacial games.
From a purely mathematical point of view, one issue that requires further attention is the proof of convergence of $f_{j}^{(\varepsilon)}\stackrel{{\scriptstyle\varepsilon\to 0}}{{\to}}\rho_{j}$, which was not established in this study. This is a necessary step in proving the rigorous compatibility between the kinetic and mesoscopic models and should be considered in future work.
For a correct modelling, the kernel $\mathfrak{Q}_{j}^{(\varepsilon)}$ required expansion of the dimensionless solution of the kinetic equation (3) in the parameter $\varepsilon$ up to order $\varepsilon^{2}$, unlike the case of the derivation of the Navier-Stokes formula from the Boltzmann equation with a collisional kernel [BGL91] or in the derivation of the Keller-Segel model from kinetic models for chemotaxis [OH02, CMPS04]. The order 2 equation, studied in Subsubsec. 3.2.3, would be trivial if there were no class-transition in the model. The fact that $\mathcal{Q}^{(T)}_{j}$ is not conservative seems to be related to the necessity of the order $\varepsilon^{2}$ expansion. Furthermore, the effect of class transition is smaller than both the diffusion (order 0) and the drift (order 1), a fact that deserves further exploration; see also [BCST26] for a recent example involving non-conservative kernels for which expansion up to order $\varepsilon^{2}$ is required to calculate formal macroscopic limits.
This paper has only begun to explore the link between kinetic models, mathematical epidemiology and, more broadly, evolutionary game theory. However, there is significant potential for further exploration of these interconnections!